# Noise Matters: Optimizing Matching Noise for Diffusion Classifiers

### v1

@@ 1 Introduction
Pretrained visual-language models (VLMs) learn the alignment of text and image from the text-image pairs. Thanks to the large-scale and in-the-wild training data, these discriminative VLMs gain some generalization capability and can even achieve zero-shot visual classification, e.g., CLIP Radford et al. (2021). When encountering some rare or customized categories, some works like prompt optimization Zhou et al. (2022b; a) can also adapt CLIP with a few-shot training set efficiently. However, the CLIP models are criticized for an inherent bag-of-words problem Yuksekgonul et al. (2022); Koishigarina et al. (2025). This problem leads to spurious bias, harming compositional inference and counterfactual reasoning. Thus, recent works Clark & Jaini (2023); Li et al. (2023); Qi et al. (2024b); Chen et al. (2023a); Yue et al. (2024) try to leverage the pretrained diffusion models for classification.
@@
The typical implementation framework is Diffusion Classifier (DC) (Clark & Jaini, 2023; Li et al., 2023). As shown in Figure 1(a), for a given clean image $x_{0}$ and a specific timestep $t$, DC randomly samples a noise $\epsilon$ and follows the forward process of diffusion (Ho et al., 2020) to get the noisy image $x_{t}$. Then input the $x_{t}$ into the diffusion denoising network $\epsilon_{\theta}$ conditioned on $K$ different categories $\{c_{i}\}_{i=1}^{K}$ respectively and get $K$ noise predictions $\{\hat{\epsilon}_{\theta}(x_{t},c_{i},t)\}_{i=1}^{K}$. The category whose corresponding noise prediction has the smallest distance to the ground truth noise $\epsilon$ is the final predicted category, i.e.,
@@
Due to the randomness of sampled noise in DC, we can sample different noises (e.g., $\epsilon_{1},\epsilon_{2}$) for the same image $x_{0}$. Unfortunately, an inherent and notorious weakness of existing DC is: Noise Instability. As shown in Figure 1(b), different sampled noises may lead to varied performance. To mitigate this instability, all prior works Clark & Jaini (2023); Li et al. (2023); Yue et al. (2024) try to sample multiple noises and calculate the expectation of the distance between them and the predicted noises for classification, i.e.,
@@
1) Frequency Matching. According to the Frequency Bias Theory Lin et al. (2022); Wang et al. (2023), given a dataset, the category-related signals are mainly in a specific frequency range. For example (c.f., Figure 1(c)), STL-10 Coates et al. (2011) contains categories “car” and “airplane”. These categories can be distinguished by their overall shapes or structures, which mainly belong to low-frequency signals. Similarly, Describable Textures (DTD) Cimpoi et al. (2014) has categories like “banded” and “dotted”. They are mainly distinguished by high-frequency signals (e.g., mutated texture). This intuitively leads to the first principle: a good noise should destroy the dataset-specific frequency signals that are related to the categories.
@@ 2 Related work
Diffusion Models (DM) for Image Classification. As a state-of-the-art generative model, DM Ho et al. (2020); Song et al. (2020) shows remarkable visual-language modeling capacity. In that case, recent studies try to unleash its discrimination potential for image classification. Studies Clark & Jaini (2023); Li et al. (2023); Chen et al. (2023a); Qi et al. (2024b) try to leverage the vision-language knowledge of pretrained text-to-image diffusion models for zero-shot classification, while Yue et.al.Yue et al. (2024) adopts the prompt optimization techniques into DM for few-shot classification. In this paper, we dive into the classification task for a new few-shot learning method of diffusion classifier.
DM for Perception Tasks. Beyond image classification, DM can also be used in more challenging perception tasks. From a framework perspective, Chen et.al. Chen et al. (2023b) leverage the DM to generate the detection bounding boxes in a refinement manner. From the pretrained knowledge perspective, studies apply the knowledge of DM in image segmentation Karazija et al. (2023); Xu et al. (2023) and depth estimation Ke et al. (2024).
Noise Instability in DM. In the image generation field, there is a widely discussed phenomenon, i.e., the start point of the denoising process matters a lot to the final generation quality. To relieve this problem, recent studies can mainly be divided into three types. 1) Finding a better sampling distribution instead of the Gaussian distribution by being supervised by a third-party model Zhou et al. (2024); Chen et al. (2024). 2) Leveraging prior knowledge Samuel et al. (2024); Rout et al. (2024); Li et al. (2024); Wang et al. (2024); Guo et al. (2024); Bai et al. (2024); Qi et al. (2024a); Bai et al. (2025) to directly refine the noise or the denoising path. 3) Directly optimizing the image-noise matching mechanism during the training process Shifeng et al. ; Liu et al. (2022). However, we find that the noise instability also exists in the image classification task, i.e., different noise leads to different classification performance. In this paper, we analyze the role of the noise and propose an optimized method to mitigate this problem.
@@ 3.1 Preliminaries
Diffusion Classifier (DC). Diffusion model (DM) Ho et al. (2020) trains a denoising network $\epsilon_{\theta}$ that can reconstruct the noisy image into the original image. To be specific, given an original image $x_{0}$, its corresponding category text $c$, and a timestep $t$, DM first samples a noise $\epsilon$ from the Gaussian distribution, then adds it to $x_{0}$ to get the noisy image $x_{t}$ by the following forward process.
@@
Prompt-Optimization of DC. The pertaining dataset of diffusion models may have a distribution gap with the downstream classification datasets. Thus, it is hard to use a few natural words to accurately describe a category directly. To fill this gap, existing few-shot DCs Yue et al. (2024) are mainly based on prompt optimization Zhou et al. (2022b; a). As shown in Figure 2(a), for each category, they set $n$ learnable text tokens $\{[V_{1}]$ $V_{2}][$$V_{n}]\}$ and add them into the text prompt to supplement the details of the category. Then they fix the whole diffusion model and only tune tokens on the few-shot training set based on Eq. (2).
@@ 3.2 Noise Optimization
After getting the classification logits, we found that all logit values are very close. This is because they are calculated based on the distance between predicted noises and the ground truth noise. And all the distances are very close due to the diffusion model’s property. Thus, if we directly optimize the $\epsilon$ and $U_{\theta}$ based on these logits, the optimization gradient is close to zero, making training difficult. To mitigate this problem, we use Z-score normalization Cheadle et al. (2003) on the category dimension to effectively amplify the signal difference between the logit. Specifically, we first calculate the mean $\mu$ and biased variance $\sigma$ of $\{p_{1},p_{2},...,p_{K}\}$ respectively:
@@ 4.1 Few-Shot Learning
Settings. To evaluate how NoOp can benefit the DC, we conducted few-shot learning experiments. Specifically, we followed the few-shot evaluation protocol of CLIP Radford et al. (2021), using 1, 2, 4, 8, and 16 shots for training, respectively, and deploying models in the full test sets. We evaluated three diffusion models, i.e., Stable Diffusion-v1.4, Stable Diffusion-v1.5 Rombach et al. (2022), Stable Diffusion-v2.0 Ramesh et al. (2022) across eight datasets: CIFAR-10 Krizhevsky et al. (2009), CIFAR-100 Krizhevsky et al. (2009), Flowers102 Nilsback & Zisserman (2008), DTD Cimpoi et al. (2014), OxfordPets Parkhi et al. (2012), EuroSAT Helber et al. (2019), STL-10 Coates et al. (2011) and FGVCAircraft Maji et al. (2013). We compared our NoOp with the ensembling methods Li et al. (2023); Clark & Jaini (2023) (i.e., sampling 5 noises for each image and consequently taking 5 times the computation for inference). For fair comparisons, we fixed the timestep $t=500$  [^1] The results of other timesteps and ensembling are in the appendix. We used the Adam optimizer Kingma & Ba (2014) with a $1e^{-2}$ and $1e^{-3}$ learning rates for the learnable noise the Meta-Network respectively. After training 20 epochs, we reported the top-1 accuracies. Results are averaged on three random seeds.
@@ 4.2 Compare with Prompt-Optimization DC
Settings. To explore the difference between noise optimization and prompt optimization. We compared the performance of a prompt-optimization based DC method  [^2] The prompt-based optimization DC is introduced in Sec. 3.1, i.e., TiF Learner Yue et al. (2024), our NoOp, and the combination of both on FGVCAircraft Maji et al. (2013) and ISIC-2019 Vinyals et al. (2016). For fairness, we used Stable Diffusion-v2.0 Ramesh et al. (2022) as the classifier. All other settings are the same as Sec. 4.1.
@@ 4.4 Ablation Study
Settings. To evaluate the effect of parameterized noise optimization (for frequency matching) and Meta-Network optimization (for spatial matching). We ablated the effectiveness of each component on 16-shot OxfordPets Parkhi et al. (2012). We used Stable Diffusion-v2.0 Ramesh et al. (2022) as the classifier and reported the top-1 accuracy and training loss across training epochs. All other settings are the same as Sec. 4.1.
@@ 4.5 Extend to Flow-based Diffusion Models
Settings. To evaluate the generalization and robustness of NoOp, we transferred it to flow-based diffusion models Lipman et al. (2022); Liu et al. (2022). We demonstrated the few-shot results of Stable Diffusion-v1.5 Rombach et al. (2022) (SD-v1.5) and the Rectified Flow Liu et al. (2022) (ReFlow, fine-tuned from SD-v1.5) on CIFAR-10 Krizhevsky et al. (2009) and OxfordPets Parkhi et al. (2012). All other settings are the same as Sec. 4.1.
@@
Results. As shown in Table 3, we have two observations: 1) The performance of ReFlow is better than SD-v1.5, although they have the same training set and architecture. This suggests that the flow-based diffusion model is also superior in the discriminative task, not just limited to the generative domain. 2) Our NoOp can not only improve the DDPM-based Ho et al. (2020) SD-v1.5 but also the flow-based Liu et al. (2022); Lipman et al. (2022) ReFlow. This indicates that our analysis of the noise role and optimization of noise is a general framework for the diffusion family.
@@
Results. As shown in Table 4, our training time is quite smaller than the inference time. Even for the 16-shot training, the total time is only around 24% while the accuracy gains 3.19% compared with the 5-times ensembling method. This demonstrates that our method can significantly improve the overall efficiency. Additional formal cost analysis is in the Appendix.
@@ 4.7 Role of Noise Validation
Settings. In this paper, we argue that the good noise in DC is trying to destroy the category-related signals. To validate this assumption, we compared two sets of noisy images. 1) Noisy images by adding random noise. 2) Noisy images by adding optimized noise (from NoOp). For fairness, we fixed $t=500$ to get the noisy images. We tested on 4 datasets: OxfordPets Parkhi et al. (2012), DTD Cimpoi et al. (2014), EuroSAT Helber et al. (2019) and Flowers102 Nilsback & Zisserman (2008), and then reported the averaged CLIP Score Hessel et al. (2021). The CLIP Score can reflect how much category-related signals are destroyed. A lower CLIP Score indicates that the noise destroys more category-related signals. we used Stable Diffusion-v2.0 Rombach et al. (2022) as the classifier and ViT-B/32 CLIP for CLIP Score calculation. All other settings are the same as Sec. 4.1.
@@ 4.8 Frequency Matching Validation
Settings. To validate our proposed frequency matching principle, we selected two representative datasets for noise optimization: 1) CIFAR-10 Krizhevsky et al. (2009), a coarse-grained dataset where the categories are mainly distinguished by low-frequency signals (object shape and structure). 2) Describable Textures (DTDCimpoi et al. (2014), a texture dataset where the categories are mainly distinguished by high-frequency signals (mutated texture). We randomly sampled one noise and then directly optimized it (w/o optimizing the Meta-Network) on these two different datasets, respectively. During the training process, we recorded the frequency change of this noise by calculating the high-frequency signal ratio of the noise. The method to get the high-frequency signal ratio is based on the 2D Fourier transform (c.f., appendix). A higher high-frequency signal ratio means this noise is a relatively high-frequency noise.
@@ 5 Conclusion
In this paper, we revealed the noise instability in the diffusion classifier and analyzed the role of noise. Then we proposed two principles as guidelines for designing the noise optimizing framework, i.e., NoOp, to refine the random noise into a stable matching noise. Extensive experiments show that our NoOp is an effective few-shot learner. It can not only learn generalized knowledge for cross-dataset recognition and be compatible with the flow matching models. Moreover, we conducted two empirical experiments to support our proposed principles. As a highlight, we find that our NoOp is orthogonal to existing optimization methods like prompt tuning. This shows that noise optimization has a unique physical meaning and effect. In the future, we are going to: 1) Extend the noise optimization into more general diffusion-based perception tasks, such as image object detection, image segmentation, or even video-domain tasks. 2) Explore whether the current unified model (e.g., auto-regressive model plus diffusion model) also has the noise instability problem, and whether we can further mitigate this problem. 3) Extend the noise optimization into flow matching, i.e., velocity optimization or sequenced velocities optimization.

### v2

@@ 1 Introduction
Pretrained visual-language models (VLMs) learn the alignment of text and image from the text-image pairs. Thanks to the large-scale and in-the-wild training data, these discriminative VLMs gain some generalization capability and can even achieve zero-shot visual classification, e.g., CLIP radford2021learning. When encountering some rare or customized categories, some works like prompt optimization zhou2022learning; zhou2022conditional can also adapt CLIP with a few-shot training set efficiently. However, the CLIP models are criticized for an inherent bag-of-words problem yuksekgonul2022and; koishigarina2025clip. This problem leads to spurious bias, harming compositional inference and counterfactual reasoning. Thus, recent works clark2023text; li2023your; qi2024simple; chen2023robust; yue2024few try to leverage the pretrained diffusion models for classification.00footnotetext: Long Chen is the corresponding author.
@@
The typical implementation framework is Diffusion Classifier (DC) (clark2023text; li2023your). As shown in Figure 1(a), for a given clean image $x_{0}$ and a specific timestep $t$, DC randomly samples a noise $\epsilon$ and follows the forward process of diffusion (ho2020denoising) to get the noisy image $x_{t}$. Then input the $x_{t}$ into the diffusion denoising network $\epsilon_{\theta}$ conditioned on $K$ different categories $\{c_{i}\}_{i=1}^{K}$ respectively and get $K$ noise predictions $\{\hat{\epsilon}_{\theta}(x_{t},c_{i},t)\}_{i=1}^{K}$. The category whose corresponding noise prediction has the smallest distance to the ground truth noise $\epsilon$ is the final predicted category, i.e.,
@@
Due to the randomness of sampled noise in DC, we can sample different noises (e.g., $\epsilon_{1},\epsilon_{2}$) for the same image $x_{0}$. Unfortunately, an inherent and notorious weakness of existing DC is: Noise Instability. As shown in Figure 1(b), different sampled noises may lead to varied performance. To mitigate this instability, all prior works clark2023text; li2023your; yue2024few try to sample multiple noises and calculate the expectation of the distance between them and the predicted noises for classification, i.e.,
@@
1) Frequency Matching. According to the Frequency Bias Theory lin2022investigating; wang2023neural, given a dataset, the category-related signals are mainly in a specific frequency range. For example (c.f., Figure 1(c)), STL-10 coates2011analysis contains categories “car” and “airplane”. These categories can be distinguished by their overall shapes or structures, which mainly belong to low-frequency signals. Similarly, Describable Textures (DTD) cimpoi2014describing has categories like “banded” and “dotted”. They are mainly distinguished by high-frequency signals (e.g., mutated texture). This intuitively leads to the first principle: a good noise should destroy the dataset-specific frequency signals that are related to the categories.
@@ 2 Related work
Diffusion Models (DM) for Image Classification. As a state-of-the-art generative model, DM ho2020denoising; song2020score shows remarkable visual-language modeling capacity. In that case, recent studies try to unleash its discrimination potential for image classification. Studies clark2023text; li2023your; chen2023robust; qi2024simple try to leverage the vision-language knowledge of pretrained text-to-image diffusion models for zero-shot classification, while Yue et.al.yue2024few adopts the prompt optimization techniques into DM for few-shot classification. In this paper, we dive into the classification task for a new few-shot learning method of diffusion classifier.
DM for Perception Tasks. Beyond image classification, DM can also be used in more challenging perception tasks. From a framework perspective, Chen et.al. chen2023diffusiondet leverage the DM to generate the detection bounding boxes in a refinement manner. From the pretrained knowledge perspective, studies apply the knowledge of DM in image segmentation karazija2023diffusion; xu2023open and depth estimation ke2024repurposing.
Noise Instability in DM. In the image generation field, there is a widely discussed phenomenon, i.e., the start point of the denoising process matters a lot to the final generation quality. To relieve this problem, recent studies can mainly be divided into three types. 1) Finding a better sampling distribution instead of the Gaussian distribution by being supervised by a third-party model zhou2024golden; chen2024find. 2) Leveraging prior knowledge samuel2024generating; rout2024semantic; li2024enhancing; wang2024silent; guo2024initno; bai2024zigzag; qi2024not; bai2025weak to directly refine the noise or the denoising path. 3) Directly optimizing the image-noise matching mechanism during the training process shifengeasing; liu2022flow. However, we find that the noise instability also exists in the image classification task, i.e., different noise leads to different classification performance. In this paper, we analyze the role of the noise and propose an optimized method to mitigate this problem.
@@ 3.1 Preliminaries
Diffusion Classifier (DC). Diffusion model (DM) ho2020denoising trains a denoising network $\epsilon_{\theta}$ that can reconstruct the noisy image into the original image. To be specific, given an original image $x_{0}$, its corresponding category text $c$, and a timestep $t$, DM first samples a noise $\epsilon$ from the Gaussian distribution, then adds it to $x_{0}$ to get the noisy image $x_{t}$ by the following forward process.
@@
Prompt-Optimization of DC. The pertaining dataset of diffusion models may have a distribution gap with the downstream classification datasets. Thus, it is hard to use a few natural words to accurately describe a category directly. To fill this gap, existing few-shot DCs yue2024few are mainly based on prompt optimization zhou2022learning; zhou2022conditional. As shown in Figure 2(a), for each category, they set $n$ learnable text tokens $\{[V_{1}]$ $V_{2}][$$V_{n}]\}$ and add them into the text prompt to supplement the details of the category. Then they fix the whole diffusion model and only tune tokens on the few-shot training set based on Eq. (2).
@@ 3.2 Noise Optimization
After getting the classification logits, we found that all logit values are very close. This is because they are calculated based on the distance between predicted noises and the ground truth noise. And all the distances are very close due to the diffusion model’s property. Thus, if we directly optimize the $\epsilon$ and $U_{\theta}$ based on these logits, the optimization gradient is close to zero, making training difficult. To mitigate this problem, we use Z-score normalization cheadle2003analysis on the category dimension to effectively amplify the signal difference between the logit. Specifically, we first calculate the mean $\mu$ and biased variance $\sigma$ of $\{p_{1},p_{2},...,p_{K}\}$ respectively:
@@ 4.1 Few-Shot Learning
Settings. To evaluate how NoOp can benefit the DC, we conducted few-shot learning experiments. Specifically, we followed the few-shot evaluation protocol of CLIP radford2021learning, using 1, 2, 4, 8, and 16 shots for training, respectively, and deploying models in the full test sets. We evaluated three diffusion models, i.e., Stable Diffusion-v1.4, Stable Diffusion-v1.5 rombach2022high, Stable Diffusion-v2.0 ramesh2022hierarchical across eight datasets: CIFAR-10 krizhevsky2009learning, CIFAR-100 krizhevsky2009learning, Flowers102 nilsback2008automated, DTD cimpoi2014describing, OxfordPets parkhi2012cats, EuroSAT helber2019eurosat, STL-10 coates2011analysis and FGVCAircraft maji2013fine. We compared our NoOp with the ensembling methods li2023your; clark2023text (i.e., sampling 5 noises for each image and consequently taking 5 times the computation for inference). For fair comparisons, we fixed the timestep $t=500$. We used the Adam optimizer kingma2014adam with a $1e^{-2}$ and $1e^{-3}$ learning rates for the learnable noise the Meta-Network respectively. After training 20 epochs, we reported the top-1 accuracies. Results are averaged on three random seeds.
@@ 4.2 Compare with Prompt-Optimization DC
Settings. To explore the difference between noise optimization and prompt optimization. We compared the performance of a prompt-optimization based DC method  [^1] The prompt-based optimization DC is introduced in Sec. 3.1, i.e., TiF Learner yue2024few, our NoOp, and the combination of both on FGVCAircraft maji2013fine and ISIC-2019 vinyals2016matching. For fairness, we used Stable Diffusion-v2.0 ramesh2022hierarchical as the classifier. All other settings are the same as Sec. 4.1.
@@ 4.4 Ablation Study
Settings. To evaluate the effect of parameterized noise optimization (for frequency matching) and Meta-Network optimization (for spatial matching). We ablated the effectiveness of each component on 16-shot OxfordPets parkhi2012cats. We used Stable Diffusion-v2.0 ramesh2022hierarchical as the classifier and reported the top-1 accuracy and training loss across training epochs. All other settings are the same as Sec. 4.1.
@@ 4.5 Extend to Flow-based Diffusion Models
Settings. To evaluate the generalization and robustness of NoOp, we transferred it to flow-based diffusion models lipman2022flow; liu2022flow. We demonstrated the few-shot results of Stable Diffusion-v1.5 rombach2022high (SD-v1.5) and the Rectified Flow liu2022flow (ReFlow, fine-tuned from SD-v1.5) on CIFAR-10 krizhevsky2009learning and OxfordPets parkhi2012cats. All other settings are the same as Sec. 4.1.
@@
Results. As shown in Table 3, we have two observations: 1) The performance of ReFlow is better than SD-v1.5, although they have the same training set and architecture. This suggests that the flow-based diffusion model is also superior in the discriminative task, not just limited to the generative domain. 2) Our NoOp can not only improve the DDPM-based ho2020denoising SD-v1.5 but also the flow-based liu2022flow; lipman2022flow ReFlow. This indicates that our analysis of the noise role and optimization of noise is a general framework for the diffusion family.
@@ 4.6 Computational Overhead Analysis
Results. As shown in Table 4, our training time is quite smaller than the inference time. Even for the 16-shot training, the total time is only around 24% while the accuracy gains 3.19% compared with the 5-times ensembling method. This demonstrates that our method can significantly improve the overall efficiency. Additional formal cost analysis is in the Appendix E.
@@ 4.7 Role of Noise Validation
Settings. In this paper, we argue that the good noise in DC is trying to destroy the category-related signals. To validate this assumption, we compared two sets of noisy images. 1) Noisy images by adding random noise. 2) Noisy images by adding optimized noise (from NoOp). For fairness, we fixed $t=500$ to get the noisy images. We tested on 4 datasets: OxfordPets parkhi2012cats, DTD cimpoi2014describing, EuroSAT helber2019eurosat and Flowers102 nilsback2008automated, and then reported the averaged CLIP Score hessel2021clipscore. The CLIP Score can reflect how much category-related signals are destroyed. A lower CLIP Score indicates that the noise destroys more category-related signals. we used Stable Diffusion-v2.0 rombach2022high as the classifier and ViT-B/32 CLIP for CLIP Score calculation. All other settings are the same as Sec. 4.1.
@@ 4.8 Frequency Matching Validation
Settings. To validate our proposed frequency matching principle, we selected two representative datasets for noise optimization: 1) CIFAR-10 krizhevsky2009learning, a coarse-grained dataset where the categories are mainly distinguished by low-frequency signals (object shape and structure). 2) Describable Textures (DTDcimpoi2014describing, a texture dataset where the categories are mainly distinguished by high-frequency signals (mutated texture). We randomly sampled one noise and then directly optimized it (w/o optimizing the Meta-Network) on these two different datasets, respectively. During the training process, we recorded the frequency change of this noise by calculating the high-frequency signal ratio of the noise. The method to get the high-frequency signal ratio is based on the 2D Fourier transform (c.f., Appendix B). A higher high-frequency signal ratio means this noise is a relatively high-frequency noise.
@@
4.9 Time steps Ablation and Ensemble
Settings. In the main paper, we fixed the timestep at 500 to focus on the noise optimization. To evaluate whether our NoOp is consistently effective on other timesteps. We conducted a timestep ablation study. We tested $t=300,400,500,600,700$ with Stable Diffusion-v2.0 ramesh2022hierarchical as the pretrained diffusion model on 4-shot CIFAR-10 krizhevsky2009learning. Moreover, we also test the ensemble with 5 different timesteps (i.e., $t=300,400,500,600,700$) after using NoOp. The pretrained diffusion model is Stable Diffusion-v2.0 ramesh2022hierarchical and dataset is 4-shot CIFAR-10 krizhevsky2009learning. We also showed the results of zero-shot w/o ensemble, zero-shot w/ensemble, and NoOp w/o ensemble for comparison.
Results. In Figure 7(a), our NoOp can gain consistent improvements across different timesteps. This indicates that NoOp is robust and effective across timesteps. Besides, as shown in Figure 7(b), we can see that compared with the one-timestep (i.e., $t=500$) NoOp, the timesteps ensemble can improve around $12\%$ top-1 accuracy and outperforms the zero-shot ensemble $5.4\%$. This demonstrates that the ensemble strategy can also be applied for our NoOp if there are enough computational resources.
@@ 5 Conclusion
In this paper, we revealed the noise instability in the diffusion classifier and analyzed the role of noise. Then we proposed two principles as guidelines for designing the noise optimizing framework, i.e., NoOp, to refine the random noise into a stable matching noise. Extensive experiments show that our NoOp is an effective few-shot learner. It can not only learn generalized knowledge for cross-dataset recognition and be compatible with the flow matching models. Moreover, we conducted two empirical experiments to support our proposed principles. As a highlight, we find that our NoOp is orthogonal to existing optimization methods like prompt tuning. This shows that noise optimization has a unique physical meaning and effect. In the future, we are going to extend the noise optimization into flow matching models and unified models (e.g., auto-regressive model plus diffusion model).
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China Young Scholar Fund (62402408) and the Hong Kong SAR RGC Early Career Scheme (26208924).
Appendix
This appendix is organized as follows:
  • Section A introduces the backgrounds of diffusion classifiers (DC), including the basic theory, implementations, and advantages compared with Vision-language Models (VLM).
  • Section B gives the theory of the 2-D Fourier transform, i.e., the frequency analysis method for images in our frequency-matching related experiments.
  • Section C gives the justification of the NoOp’s mechanism.
  • Section D provides more experimental results. Firstly, we give a few additional few-shot learning results compared with the prompt-optimization DC in Section D.1. Secondly, we provide the ablation results of Meta-Network with different architectures in Section D.2. Finally, we give the qualitative results of our NoOp’s stability in Section D.3.
  • Section E provides the implementation details of our NoOp and the reproduction details. Besides, we also give the computation cost.
  • Section F analyzes the limitations of noise optimization DC and its societal impacts.